. A rectangular garden is 3 times as long as it is wide. If the length is decreased by 8 ft and the width is increased by 5 ft, the perimeter will be 66 ft. Find the dimensions of the original garden.
Answer by lwsshak3(11628)

A rectangular garden is 3 times as long as it is wide. If the length is decreased by 8 ft and the width is increased by 5 ft, the perimeter will be 66 ft. Find the dimensions of the original garden.
***
original dimensons:
let x=width of garden
3x=length of garden
..
new dimensions:
x+5=width
3x-8=length
..
2*(length+width)=perimeter
2(3x-8+x+5)=66
2(4x-3=66
4x-3=33
4x=36
x=9
3x=27
dimensions of the original garden:
width=9 ft
length=27 ft